### v1

# skwdro: a library for Wasserstein distributionally robust machine learning

Abstract
We present skwdro, a Python library for training robust machine learning models. The library is based on distributionally robust optimization using optimal transport distances. For ease of use, it features both scikit-learn compatible estimators for popular objectives, as well as a wrapper for PyTorch modules, enabling researchers and practitioners to use it in a wide range of models with minimal code changes. Its implementation relies on an entropic smoothing of the original robust objective in order to ensure maximal model flexibility. The library is available at https://github.com/iutzeler/skwdro.
Keywords: Distributionally robust optim., distribution shifts, entropic regularization
1 Introduction: ERM, WDRO, and entropic regularization
Training machine learning models typically relies on empirical risk minimization (ERM), which consists in minimizing the expectation of the loss of a model on the empirical distribution of training data. This approach has been questioned for its limited resilience and reliability, as machine learning systems have become widely used and deployed for prolonged periods of times. Recently, distributionally robust optimization (DRO) has emerged as a powerful paradigm towards better generalization and better resilience against distribution shifts, data heterogeneity, and adversarial attacks, see e.g., the monograph Chen et al. (2020) Blanchet et al. (2021)
The paradigm consists in minimizing the average loss over the worst distribution in a neighborhood of the empirical distribution, thus capturing any data uncertainty. A natural way to define this neighborhood is through the optimal transport distance (called Wasserstein distance, see e.g., the textbook Peyré and Cuturi (2019)), as proposed by (Esfahani and Kuhn, 2018). This framework, called Wasserstein distributionally robust optimization (WDRO), is attractive, since it combines powerful modeling capabilities, strong generalization properties, and practical robustness guarantees; see e.g., recent theoretical and practical developments (Sinha et al., 2018; Kuhn et al., 2019; Azizian et al., 2024; Blanchet et al., 2024). Unfortunately, the resulting optimization problem is usually intractable, which limits the use of WDRO in practice for training robust models.
In this work, we tackle this computational issue and provide a complete and easy-to-use library for numerical WDRO. In Section 2, we formalize the notation and sketch our approach. We present the library skWDRO in Section 3 and mention the numerical ingredients in Section 4. The online documentation completes this short article by presenting detailed features and numerical illustrations.
2 WDRO models and entropic variants
Formally, we aim at selecting a model among a family parametrized by ${\theta}\in{\Theta}$, whose error is measured by some loss function ${f}_{\theta}({\xi})$. Given ${n}$ data points $({\xi}_{i})_{{i}=1}^{n}$ (raw data or input-label pairs), the ERM problem writes
$\min_{{\theta}\in{\Theta}}\leavevmode\nobreak\ \leavevmode\nobreak\ {\frac{1}{{n}}\sum_{{i}=1}^{n}}{f}_{\theta}({\xi}_{i})=\mathbb{E}_{{\xi}\sim\operatorname{\operatorname{P}_{{n}}}}\left[{f}_{\theta}({\xi})\right]\quad\text{ where the empirical distribution is }\operatorname{\operatorname{P}_{{n}}}={\frac{1}{{n}}\sum_{{i}=1}^{n}}\delta_{{\xi}_{i}}.$ (1)
As mentioned in introduction, this approach can fail if the number of data points is too small or the data distribution at deployment differs from the training distribution. WDRO is a way to remedy these issues: it consists in minimizing the worst expectation of the loss when the probability distribution $Q$ lives in a Wasserstein ball at the empirical distribution $\operatorname{\operatorname{P}_{{n}}}$. Thus, for a robustness radius ${\rho}\geq 0$, this leads to the WDRO problem
$\min_{{\theta}\in{\Theta}}\leavevmode\nobreak\ \sup_{W_{{c}}(\operatorname{\operatorname{P}_{{n}}},\operatorname{Q})\leq{\rho}}\mathbb{E}_{{\xi}\sim\operatorname{Q}}\left[{f}_{\theta}({\xi})\right]\,,$ (2)
where $W_{{c}}(\operatorname{\operatorname{P}_{{n}}},\operatorname{Q})$ denotes the optimal transport cost between $\operatorname{\operatorname{P}_{{n}}}$ and $\operatorname{Q}$ with ground cost ${c}:{\Xi}\times{\Xi}\to\mathbb{R}_{+}$, often referred to as the Wasserstein distance. Applying the standard Lagrangian duality gives the following dual problem
$\min_{{\theta}\in{\Theta}}\leavevmode\nobreak\ \min_{{\lambda}\geq 0}\leavevmode\nobreak\ {\lambda}{\rho}+{\varepsilon}\leavevmode\nobreak\ \mathbb{E}_{{\xi}\sim\operatorname{\operatorname{P}_{{n}}}}\left[\sup_{\zeta}\leavevmode\nobreak\ \left\{{f}_{\theta}({\zeta})-{\lambda}{c}({\xi},{\zeta})\right\}\right]$ (3)
where ${\lambda}\geq 0$ is the dual variable associated to the Wasserstein distance constraint. For some specific tasks, this problem can be reformulated as a convex program (Kuhn et al., 2019; Shafieezadeh-Abadeh et al., 2019). However, this convex approach does not allow for complex models, such as neural networks. Thus, though very attractive, the robust approach usually leads to an intractable optimization problem.
To remedy this issue, we consider the following smoothed counterpart of (3)
$\min_{{\theta}\in{\Theta}}\leavevmode\nobreak\ \min_{{\lambda}\geq 0}\leavevmode\nobreak\ {\lambda}{\rho}+{\varepsilon}\leavevmode\nobreak\ \mathbb{E}_{{\xi}\sim\operatorname{\operatorname{P}_{{n}}}}\left[\log\left(\mathbb{E}_{{\zeta}\sim\mathcal{N}({\xi},\sigma^{2}I)}\left[\exp{\left(\frac{{f}_{\theta}({\zeta})-{\lambda}{c}({\xi},{\zeta})}{{\varepsilon}}\right)}\right]\right)\right]$ (4)
where the regularization strength ${\varepsilon}$ and the sampling variance $\sigma^{2}$ are the two parameters controlling the approximation. This problem interprets as the dual of the entropy regularized WDRO problem (Azizian et al., 2023; Wang et al., 2023). Interestingly, such entropic-regularized problems have been proved to retain the generalization and robustness guarantees of (3); see e.g., Azizian et al. (2024).
In this work, we leverage the practical interest of the approach: (4) can be solved using (stochastic) gradient-based methods, which thus significantly broadens the applicability and numerical tractability of distributionally robust machine learning. It is at the core of skwdro.
3 skwdro: implementation and features
skwdro is an open-source library (https://github.com/iutzeler/skwdro) for machine learning with WDRO written in Python, based on the methodology presented in the previous section. Its user interface is twofold. In the case where (2) can be reformulated as a standard convex problem, skwdro can be used directly through its scikit-learn interface (Pedregosa et al., 2011), which relies on cvxpy for optimization (Diamond and Boyd, 2016). For all other cases, skwdro can transform any PyTorch model (Paszke et al., 2019) into a robust counterpart, via the entropic smoothing (4). Overall, skwdro allows to robustify machine learning models with minimal code changes, as illustrated in the next subsections.
Figure 1: Schematic view of the main blocks of skwdro.
3.1 scikit-learn interface
In skwdro, standard models such, as logistic/linear regression, are directly available, implemented as scikit-learn estimators. They thus benefit from the same interface and the seamless integration with the rest of scikit-learn’s functionalities such as scoring, pipelining, cross-validation, etc. As an example, we show below the fitting of scikit-learn’s LinearRegression estimator and its robust counterpart with skwdro.
1from sklearn.linear_model import LinearRegression # sklearn’s regressor
2from skwdrwo.linear_models import LinearRegression as RobustLinearRegression
3X,y = ... # training data
4# $=$ ERM $=$
5erm_lin = LinearRegression()
6erm_lin.fit(X,y)
7# $=$ DRO $=$
8rob_lin = RobustLinearRegression(rho=0.1) # WDRO with radius 0.1
9rob_lin.fit(X,y)
3.2 PyTorch interface
skwdro provides an easy-to-use interface for PyTorch functionalities. In practice, it enables users to robustify PyTorch models, with a minimal change in their code. In the example below, the standard training loop (in red) is replaced by its robust counterpart (in green) using skwdro’s wrapper robustify.
1import torch as pt
2from skwdro.torch import robustify
3model = # your model as a PyTorch Module
4loss_fn = # your loss as a PyTorch Module
5wdro_loss = robustify(
6 loss_fn, model,
7 rho = 0.1, # robustness radius
8 X_init, y_init # provide some sample data for shape initialization
9) # replaces the loss of the model by the dual WDRO loss
10optimizer = # your choice of optimization method
11for epoch in range(num_epochs): # training loop
12 for X, y in train_batches:
13 optimizer.zero_grad()
14 # $=$ ERM $=$ Here is what you would usually do to train the model:
15 # loss = loss_fn(model(X), y).mean() # standard loss on batch
16 # loss.backward()
17 # optimizer.step() # optimization step
18 # $=$ DRO $=$ Here is what you would do to train the robust model:
19 rob_loss = wdro_loss(X, y).mean() # robust loss
20 rob_loss.backward()
21 optimizer.step() # optimization step
4 Numerical aspects
The generic robustification of PyTorch models through the smoothed objective (4) presents some numerical challenges, before even optimizing the model, indeed:
  • Given an implementation of the loss function ${f}_{\theta}({\xi})$, the objective (4) is not yet amenable to stochastic first-order optimization, due to the presence of a (continuous) expectation inside the logarithm. We approximate this integral by Monte-Carlo sampling, with a bound on the induced bias.
  • The presence of the exponential in (4) makes the objective sharply peaked, which results in a high variance in the gradient estimate. We mitigate this with importance sampling, shifting the samples following the gradient $\nabla_{\!{\xi}}{f}_{\theta}({\xi}_{i})$.
  • The smoothing depends on some parameters $\sigma$, $\varepsilon$ that have to be tuned. We provide an automatic calibration for them, based on the problem parameters and the following theoretical guidelines of Azizian et al. (2024).
Once the model loss is robustified, the user has two options: either use the new loss in their own training/optimization loop, or use the optimization procedure provided by the library (relying on adaptive optimizers (Cutkosky et al., 2023)).
5 Conclusion
skwdro is a publicly-available library for WDRO machine learning. Through the implementation of the whole pipeline and user interfaces, the goal of the library is to make the training of robust model easy for a wide audience. At the core of the implementation, is a careful numerical work on entropic smoothing of the WDRO objective. Numerical illustrations as well as features and implementation details are available in the online documentation at https://skwdro.readthedocs.io.
Acknowledgments and Disclosure of Funding
This work has been partially supported by MIAI@Grenoble Alpes (ANR-19-P3IA-0003)